A car travels for 90 miles at an average speed 60mph. Due to traffic it then reduces its average speed to 30mph for a further 60 miles. How long does the journey take?

Answer by mananth(16949)

A car travels for 90 miles at an average speed 60mph. Due to traffic it then reduces its average speed to 30mph for a further 60 miles. How long does the journey take?
..
Speed = 60 mph
distance = 90 miles
t=d/r
t=90/60
t=3/2 hours
= 1.5 hours
...
speed = 30mph
distance = 60 miles
..
t=d/r
t=60/30
t=2 hours.
...
Total time = 1.5 hours + 2 hours
Total time = 3.5 hours.
m.ananth@hotmail.ca